OpenStudy (anonymous):

Having trouble with this. Find the equation of the tangent. f(x)=x^2-2; a=4

OpenStudy (anonymous):

To find the equation of the tangent, you first need to find the slope of the tangent line. THhe slope fo the tangent is found using the derivative. f'(x)=2x f(4)=8. So, the slope of the tangent line is 8. Now, for the equation of the line, you need a point. By plugging 4 into your original function, you can get the tangent point to the curve. f(4)=16-2=14 so (4,14) is the tangent point. So, use y-b=m(x-a) to get your tangent line equation using (4,14) as (a,b) and m=8.

OpenStudy (anonymous):

Ohhhhhhh. I kept using 2 from f'(x) as my slope. Thank you. Clears everything up.

OpenStudy (anonymous):

That would make quite the difference. :) Glad I could help.